40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 24 16:59:31 2012, Jenne, Update, VAC, Doors on, ready to pump 
    Reply  Mon Sep 24 17:47:05 2012, Steve, Update, VAC, ready to pump 
       Reply  Tue Sep 25 17:21:01 2012, Steve, Update, VAC, pumpdown completed pd73.pngpd73b.png
          Reply  Wed Sep 26 11:03:04 2012, Steve, Update, VAC, pumpdown completed 40mpd73md1.png28datm.png
             Reply  Mon Oct 1 09:02:17 2012, Steve, Update, VAC, pumpdown reached vacuum normal pd73md6.pngpd73md6b.png
                Reply  Tue Oct 2 09:13:16 2012, Steve, Update, VAC, RGA scan at day 7 pd73md7.png
                Reply  Tue Oct 9 09:29:10 2012, Steve, Update, IOO, this is where we are 15days.png
    Reply  Mon Oct 8 22:07:46 2012, rana, Update, SUS, Doors on, ready to pump 20120921_200637.jpg
Message ID: 7432     Entry time: Mon Sep 24 16:59:31 2012     Reply to this: 7434   7507
Author: Jenne 
Type: Update 
Category: VAC 
Subject: Doors on, ready to pump 

[The 40m Family]

The access connector and all heavy doors are back on. 

Jamie put the regular viton EQ stops back on PRM, since he had to adjust the distance between the EQ stops and the PRM anyway.  Jamie also waved an IR card near the IPANG steering mirrors in ETMY, but it was not possible to take a good photo.  Jamie certifies that the beam is centered on both of those 2 optics.

I have centered IPPOS and IPANG QPDs. 

All oplevs need a little realignment, especially ETMY, which had it's lens removed (Rana has a Wall of Shame photo of this, which is why it was removed by him).  Steve will look into this tomorrow, after he starts pumping.

I have turned off all PZT high voltage supplies for in-vac PZTs: The input PZTs, the output PZTs, and the OMC PZTs (which weren't on, but I confirmed they were off).

I have also prepared the 3 low-power items for high power: MC refl's path was changed back to regular BS, AS camera was moved to its nominal position, and IPPOS has its ND filters back.  MC refl and the AS camera will need to be realigned once we're actually at high power tomorrow afternoon.

Long vent, but good work everyone.

ELOG V3.1.3-